### v2

@@ 8 Acknowledgments
$e_{Yik}=\theta_{Y0ik}^{\prime}+\theta_{Vk}\alpha_{0ik}^{\prime}+\theta_{Vk}\alpha_{1ik}^{\prime}Z_{ik}+\gamma_{1ik}^{\prime}Z_{ik}+\gamma_{2k}e_{Dik}+\gamma_{2ik}^{\prime}\left(\beta_{0k}+\beta_{1k}Z_{ik}+e_{Dik}\right)+\gamma_{3k}e_{Dik}Z_{ik}+\gamma_{3ik}^{\prime}\left(\beta_{0k}Z_{ik}+\beta_{1k}Z_{ik}+e_{Dik}Z_{ik}\right)$.
It is obvious that $E\left(e_{Yik}\mid S_{i}=k\right)=E\left(e_{Yik}Z_{ik}\mid S_{i}=k\right)=0$ if $\operatorname{Cov}\left(\gamma_{3ik}^{\prime},e_{Dik}\mid S_{i}=k\right)=\operatorname{Cov}\left(\gamma_{2ik}^{\prime},e_{Dik}\mid S_{i}=k\right)=0$. These conditions hold under Assumption 2.
@@
Step 4. To generate a posttreatment confounder $V_{ik}(z)$ for $z=0,1$, we first generate $\phi_{0ik}$ and $\phi_{1ik}$ each as a function of $U_{ik}$ and $X_{ik}$ with different site-specific intercepts $t_{0k}$ and $t_{1k}$, respectively. Here $\phi_{0ik}=P(V_{ik}=1\mid S_{i}=k,Z_{ik}=0,U_{ik},X_{ik})$ and $\phi_{1ik}=P(V_{ik}=1\mid S_{i}=k,Z_{ik}=1,U_{ik},X_{ik})$. To simplify, the slopes for $U_{ik}$ and $X_{ik}$ are constant across the sites and do not differ for $\phi_{0ik}$ and $\phi_{1ik}$.
@@
[…]
We force the values of $\phi_{0ik}$ and $\phi_{1ik}$ to be within $[0,1]$. The corresponding key parameters in our theoretical model are as follows:
@@
[…]
where $\alpha_{0ik}=\phi_{0ik}$ and $\alpha_{1ik}=\phi_{1ik}-\phi_{0ik}$.
Finally, we generate the binary $V_{ik}(0)$ and $V_{ik}(1)$ by taking random draws from their respective Bernoulli distributions with parameter values $\phi_{0ik}$ and $\phi_{1ik}$, and $V_{ik}=Z_{ik}V_{ik}(1)+(1-Z_{ik})V_{ik}(0)$.
Step 5. To generate the Phase II treatment assignment $D_{ik}$, we first generate $\pi_{zik}$ as a function of $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$, for $z=0,1$, where $\pi_{0ik}$ and $\pi_{1ik}$ have different site-specific intercepts $s_{0k}$ and $s_{1k}$, respectively. Here $\pi_{0ik}=Pr(D_{ik}=1\mid S_{i}=k,Z_{ik}=0,U_{ik},X_{ik},V_{ik}(0))$ and $\pi_{1ik}=Pr(D_{ik}=1\mid S_{i}=k,Z_{ik}=1,U_{ik},X_{ik},V_{ik}(1))$. Given $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$, the Phase II treatment assignment $D_{ik}(z)$ still follows a random process given the stochasticity in the real world. The slopes for $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$ are each a function of $z$. To simplify, each of these slopes is constant across the sites.
[…]
where $s_{0k}\sim N(0,0.08^{2})$ and $s_{1k}\sim N(0,0.1^{2})$ are independent.
We force the values of $\pi_{0ik}$ and $\pi_{1ik}$ to be within $[0,1]$. Using this model, the corresponding key parameters in our theoretical model are as follows:
$D_{ik}(z)=$ $\beta_{D0k}+\beta_{X0k}(X_{ik}-\bar{X}_{k})+\beta_{U0k}(U_{ik}-\bar{U}_{k})+\beta_{V0k}V_{ik}(0)+\beta_{D1k}z$
$+(\beta_{X1k}-\beta_{X0k})(X_{ik}-\bar{X}_{k})z+(\beta_{U1k}-\beta_{U0k})(U_{ik}-\bar{U}_{k})z$
$+\{\beta_{V1k}V_{ik}(1)-\beta_{V0k}V_{ik}(0)\}z+e_{z.ik}^{(D)},$
[…]
Finally, we generate the binary $D_{ik}$ by taking a random draw from a Bernoulli distribution with parameter value $\pi_{ik}=Z_{ik}\pi_{1ik}+(1-Z_{ik})\pi_{0ik}$.
Step 6. Lastly, we generate a continuous $Y_{ik}(z,d)$ for $z=0,1$ and $d=0,1$, each as a function of $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$. To simplify, the slopes for $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$ are constant across sites. Consistent with the theoretical model, the slopes for $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$ do not depend on $z$ or $d$.
@@
$ATE=\gamma_{1}+\gamma_{2}+\gamma_{3}+\alpha_{1}\theta_{V}=10+15-5+(-0.1)\times(-15)=21.5.$

### v3

@@ 8 Acknowledgments
Since $\epsilon_{k}=\left(\gamma_{1k}-\gamma_{1}\right)+\left(\gamma_{2k}-\gamma_{2}\right)\beta_{1k}+\left(\gamma_{3k}-\gamma_{3}\right)\beta_{2k}+\left(\theta_{Vk}-\theta_{V}\right)\alpha_{1k}$, and therefore,
$\mathrm{E}(\epsilon_{k})$ $=\mathrm{E}\left(\gamma_{1k}-\gamma_{1}\right)+\mathrm{E}\{\left(\gamma_{2k}-\gamma_{2}\right)\beta_{1k}\}+\mathrm{E}\{\left(\gamma_{3k}-\gamma_{3}\right)\beta_{2k}\}+\mathrm{E}\{\left(\theta_{Vk}-\theta_{V}\right)\alpha_{1k}\}$
$=0+\mathrm{E}\left(\gamma_{2k}-\gamma_{2}\right)\mathrm{E}(\beta_{1k})+\mathrm{E}\left(\gamma_{3k}-\gamma_{3}\right)\mathrm{E}(\beta_{2k})+\mathrm{E}\left(\theta_{Vk}-\theta_{V}\right)\mathrm{E}(\alpha_{1k})$
$=0,$
where the second identity holds under Assumption 3.
$\mathrm{Cov}(\epsilon_{k},\beta_{1k})$ $=\mathrm{Cov}(\gamma_{1k}-\gamma_{1},\beta_{1k})+\mathrm{Cov}\{(\gamma_{2k}-\gamma_{2})\beta_{1k},\beta_{1k}\}+\mathrm{Cov}\{(\gamma_{3k}-\gamma_{3})\beta_{2k},\beta_{1k}\}$
$+\mathrm{Cov}\{(\theta_{Vk}-\theta_{V})\alpha_{1k},\beta_{1k}\}.$
In the above expression, $\mathrm{Cov}(\gamma_{1k}-\gamma_{1},\beta_{1k})=0$ given the independence between $\gamma_{1k}$ and $\beta_{1k}$ under Assumption 3. The term $\mathrm{Cov}\{(\gamma_{2k}-\gamma_{2})\beta_{1k},\beta_{1k}\}=0$ given the independence between $\gamma_{2k}$ and $\beta_{1k}$ under Assumption 3. The term $\mathrm{Cov}\{(\gamma_{3k}-\gamma_{3})\beta_{2k},\beta_{1k}\}=0$ given the joint independence between $(\beta_{1k},\beta_{2k})$ with $\gamma_{3k}$ under Assumption 3. The term $\mathrm{Cov}\{(\theta_{Vk}-\theta_{V})\alpha_{1k},\beta_{1k}\}=0$ given the joint independence between $(\alpha_{1k},\beta_{1k})$ with $\theta_{Vk}$ under Assumption 3.
Similarly, one can show that $\mathrm{Cov}(\epsilon_{k},\beta_{2k})$ and $\mathrm{Cov}(\epsilon_{k},\alpha_{1k})$ are both zero under Assumption 3.
@@
Step 4. We generate the Phase I outcome $V_{ik}(z)$ for $z=0,1$, each as a function of school-mean-centered $U_{ik}$ and $X_{ik}$ with different site-specific intercepts. To simplify, the slopes for $U_{ik}$ and $X_{ik}$ are constant across the sites, and to satisfy the within-site zero covariance assumption (Assumption 2), they do not differ for $V_{ik}(0)$ and $V_{ik}(1)$.
@@
where $t_{0k}\sim N(0,8^{2})$ and $t_{1k}\sim N(0,6^{2})$ are independent.
The corresponding parameters in our theoretical model are as follows:
@@
$\alpha_{1ik}$ $=5+t_{1k}.$
The observed Phase I outcome is $V_{ik}=Z_{ik}V_{ik}(1)+(1-Z_{ik})V_{ik}(0)$.
Step 5. To generate the Phase II treatment assignment $D_{ik}(z)$ for $z=0,1$, we specify a pair of logit models as follows:
$D_{ik}(0)$ $=\mathbbm{1}\left\{-(X_{ik}-\bar{X}_{k})-(U_{ik}-\bar{U}_{k})-0.1V_{ik}(0)+s_{0k}-u_{ik}^{(0)}\geq 0\right\},$
$D_{ik}(1)$ $=\mathbbm{1}\left\{X_{ik}-\bar{X}_{k}+U_{ik}-\bar{U}_{k}+0.05V_{ik}(1)+s_{1k}-u_{ik}^{(1)}\geq 0\right\},$
where $s_{0k}\sim N(0,1)$ and $s_{1k}\sim N(0,1)$ are independent and each i.i.d. across sites, $u_{ik}^{(0)}$ and $u_{ik}^{(1)}$ are independent i.i.d. Logistic(0, 1) random variables.
The observed Phase II treatment is $D_{ik}=Z_{ik}D_{ik}(1)+(1-Z_{ik})D_{ik}(0)$.
Step 6. Lastly, we generate a continuous $Y_{ik}(z,d)$ for $z=0,1$ and $d=0,1$, each as a function of $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$. To simplify, the slopes for $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$ are constant across sites. To satisfy the within-site zero covariance assumption (Assumption 2), the slopes for $U_{ik}$, $X_{ik}$, and $V_{ik}(z)$ do not depend on $z$ or $d$.
@@
$ATE=\gamma_{1}+\gamma_{2}+\gamma_{3}+\alpha_{1}\theta_{V}=10+15-5+5\times 0.2=21.$
In the second set of simulations, in order to examine the consequence of violating the within-site zero covariance assumption (Assumption 2), we make $\gamma_{1ik}$, $\gamma_{2ik}$, and $\gamma_{3ik}$ each be a function of $X_{ik}$ and $U_{ik}$ in the model for $Y_{ik}(z,d)$. Specifically,
[…]
$+\gamma_{1k}z+\gamma_{2k}d+\gamma_{3k}zd+e_{zd.ik}^{(Y)}.$
[…]
$\gamma_{3ik}$ $=-5+g_{zdk}+5(U_{ik}-\bar{U}_{ik}).$
Note that $\beta_{1ik}$ and $\beta_{2ik}$ are already each a function of $X_{ik}$ and $U_{ik}$ in the model for $D_{ik}(z)$. Therefore, $\mathrm{Cov}(\gamma_{2ik},\beta_{1ik})\neq 0$, $\mathrm{Cov}(\gamma_{3ik},\beta_{2ik})\neq 0$, and the within-site zero covariance assumption (Assumption 2) is violated.
In the third set of simulations, we examine the consequence of violating the between-site independence assumption (Assumption 3) by replacing $g_{zk}$ in the $Y$ model by $\frac{1}{3}t_{1k}$, so that $\alpha_{1k}=5+t_{1k}$ and $\gamma_{1k}=10+\frac{1}{3}t_{1k}$ are correlated.
@@
S5 Estimation Performance When the Assumptions are Violated
Table S1: Estimation Performance (Assumption 2 Violated)
$K=25$ $K=100$
$n_{k}$ 30 100 1,000 5,000 30 100 1,000 5,000
Bias
Naive 2.75 2.71 2.75 2.76 2.69 2.70 2.75 2.71
Naive-adj 2.05 2.04 2.10 2.11 1.93 2.06 2.11 2.06
IPTW -0.99 -0.31 0.51 0.65 -1.16 -0.30 0.45 0.53
MS2T-IV 0.04 0.09 0.17 0.12 0.05 0.11 0.18 0.16
MS2T-IV-adj -0.01 0.05 0.15 0.12 0.01 0.05 0.16 0.15
Empirical Variance
Naive 7.09 3.50 1.18 1.09 1.92 0.74 0.31 0.28
Naive-adj 6.46 3.02 0.91 0.87 1.88 0.65 0.24 0.22
IPTW 29.67 9.94 5.04 4.42 7.86 2.43 1.30 1.26
MS2T-IV 11.43 5.27 1.82 1.68 2.65 1.40 0.47 0.32
MS2T-IV-adj 10.47 4.37 1.73 1.64 2.47 1.22 0.44 0.31
MSE
Naive 14.65 10.82 8.76 8.72 9.16 8.03 7.89 7.63
Naive-adj 10.68 7.17 5.33 5.32 5.62 4.90 4.69 4.47
IPTW 30.65 10.04 5.30 4.84 9.20 2.52 1.50 1.54
MS2T-IV 11.44 5.28 1.85 1.70 2.65 1.41 0.50 0.35
MS2T-IV-adj 10.47 4.37 1.75 1.66 2.47 1.23 0.46 0.33
Table S2: Estimation Performance (Assumption 3 Violated)
$K=25$ $K=100$
$n_{k}$ 30 100 1,000 5,000 30 100 1,000 5,000
Bias
Naive 1.80 1.73 1.79 1.83 1.75 1.74 1.78 1.74
Naive-adj 1.36 1.30 1.37 1.41 1.27 1.33 1.37 1.33
IPTW -2.62 -1.52 -0.51 -0.34 -2.83 -1.50 -0.59 -0.50
MS2T-IV -0.01 0.00 0.03 0.03 -0.02 -0.04 0.03 0.00
MS2T-IV-adj -0.03 -0.01 0.01 0.03 -0.04 -0.09 0.01 0.00
Empirical Variance
Naive 5.34 2.67 1.06 0.99 1.45 0.58 0.26 0.26
Naive-adj 4.92 2.40 0.88 0.86 1.41 0.52 0.22 0.22
IPTW 25.33 8.82 4.80 4.17 6.67 2.15 1.22 1.21
MS2T-IV 7.52 3.40 1.26 1.23 1.81 0.82 0.30 0.24
MS2T-IV-adj 6.80 2.99 1.18 1.21 1.67 0.73 0.28 0.23
MSE
Naive 8.57 5.68 4.25 4.33 4.50 3.60 3.42 3.29
Naive-adj 6.78 4.10 2.75 2.84 3.02 2.30 2.08 1.97
IPTW 32.20 11.12 5.06 4.28 14.65 4.41 1.57 1.46
MS2T-IV 7.52 3.40 1.27 1.23 1.81 0.83 0.30 0.24
MS2T-IV-adj 6.81 2.99 1.18 1.21 1.67 0.74 0.28 0.23